OpenStudy (anonymous):

What are the directrices of the ellipse given by the equation x^2/36+y^2/100=1?

OpenStudy (anonymous):

i think thats it anyway

OpenStudy (anonymous):

tell me if im right

OpenStudy (anonymous):

My answers I have to chose from are x=pm 5/8 y=pm 5/8 x=pm 25/2 y=pm 25/2